OpenStudy (anonymous):

Help? @johnweldon1993

OpenStudy (anonymous):

As part of a science project on plant food, Alicia and Hyun each measured a plant’s growth over several weeks. Their data is shown below. Alicia’s Plant’s Growth Week 1 2 3 4 5 Height (in cm) 3.05 6.1 9.15 12.2 15.25 Which of the following statements best reflects Alicia’s and Hyun’s data? A. Alicia’s plant showed a greater rate of growth than Hyun’s plant did. B. Hyun’s plant showed a greater rate of growth than Alicia’s plant did. C. Alicia’s plant and Hyun’s plant showed the same rate of growth. D. The rates of growth for both Alicia’s plant and Hyun’s plant changed over time. @jim_thompson5910 @johnweldon1993

OpenStudy (anonymous):

@johnweldon1993 @mathmale @Hero @ganeshie8 @tHe_FiZiCx99 @iPwnBunnies @mathmale @marissalovescats @jim_thompson5910 @nincompoop

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

@phi @Compassionate @mathstudent55 @wio @kropot72 @hba A?

hero (hero):

Explain what you did to arrive at A as a possible choice.

OpenStudy (anonymous):

It says in 5 weeks Alicia's Plant grew 15.25 and Hyun's 15.

OpenStudy (anonymous):

Actually Hyun's doesn't look exactly 15, so I'm going to change that to C. @Hero

OpenStudy (mathstudent55):

Can you use the graph of Hyun's experiment and make a table similar to the table you have for Alicia? Read the points off the graph for weeks 1, 2, 3, 4, 5.

hero (hero):

@CrazyCountryGirl, we talked about this the other day. Rate of Change is just a phrase for slope.

OpenStudy (anonymous):

week 1= 3 week 2= 5 week 3= 8 week 4= 11 week 5= 14

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

It looks like Alicia's changes are bigger....

OpenStudy (mathstudent55):

That's approximately correct. Now do this. Place Hyun's results below Alicia's results and compare them.

OpenStudy (anonymous):

A

OpenStudy (mathstudent55):

Week 1 2 3 4 5 Height (in cm) Alicia 3.05 6.1 9.15 12.2 15.25 Height (in cm) Hyun 3 5 8 11 14

OpenStudy (anonymous):

A?

OpenStudy (mathstudent55):

We can come to a few conclusions: Alicial plant growth is greater than Hyun's. Hyun's results are definitely linear because the graph shows a line. Alicia's results are also linear because 1 * 3.05 = 3.05 2 * 3.05 = 6.1 3 * 3.05 = 9.15 etc. With the above in mind, the answer is clearly A.